# Resolving power: A general approach to compare the distinguishing ability of threshold-free evaluation metrics

### v2

@@ 6 Empirical sampling models
The last row uses the signal curve to map AUPRC to the AUROC scale. Recall that metric resolution, $\kappa$, is just the width of the 95 percent confidence interval in AUROC units. The AUROC has considerably better discrimination with a resolving power that is over 70 percent greater than the AUPRC. The absolute difference in confidence interval widths is about 0.011 AUROC units, which could be substantial relative to the often small improvements achieved during hyperparameter tuning.
The robustness of these results with respect to the additive improvement assumption could be checked with a sensitivity analysis that explores other paths towards class score separation. An extreme alternative would be to resolve errors starting with the highest risk scores first and then moving down the rankings. This would yield an immediate strong response for the AUPRC while the AUROC would maintain its constant rate of improvement. An opposite extreme would be to improve all the lowest ranked cases first. There are also myriad intermediate strategies. One could use a mechanism similar to Mazzanti [14], with improvements initially evenly distributed across the score distribution but then gradually becoming concentrated in the early retrieval area.
@@ 7 Conclusion
Evaluation metrics form the contours of a model’s performance topography, so choosing the right metric is essential for successful navigation of this space. Resolving power is a framework for comparing threshold-free metrics and selecting the best for a particular problem. Central to the method is the specification of a class score sampling model that is used to both manipulate model quality and probe sampling variability. The quality dimension, which serves as the standard for comparison, is an ordered sequence of score distributions with decreasing overlap. This paper uses simple additive shifts to separate the class scores. The importance of this choice can be explored with sensitivity analyses that test robustness to alternate sequences of distributions. Signal curves show how evaluation metrics respond to changes in classifier quality. Metric error variance is found with random draws from the sampling model. Resolving power is classifier-agnostic since it operates on risk scores that are downstream of a classification model. Binormal model simulation results provide general rules-of-thumb for when the AUROC will have stronger resolving power than the AUPRC. The empirical method allows researchers to use their data and an initial classifier of their choice to construct a sampling model for estimating resolving power.
@@
All code and simulation data used in the paper is available on GitHub at https://github.com/colinbeam/resolving_power
@@ Appendix A Linear approximation method for resolving power
Figure 8: Signal curve plotting the AUROC references scale on the y-axis. A linear approximation of the signal function at the evaluation point is shown in red.
@@
Inequality A.1 makes transparent the signal to noise comparison. The left side of the inequality is the relative signal of the two metrics while the right side is the relative noise. If the inequality holds it means that the signal of the AUROC overwhelms its noise, giving it relatively greater resolving power.

### v3

@@ 6 Empirical sampling models
The last row uses the signal curve to map AUPRC to the AUROC scale. Recall that metric resolution, $\kappa$, is just the width of the 95 percent confidence interval in AUROC units. The AUROC has considerably better discrimination with a resolving power that is over 70 percent greater than the AUPRC. The absolute difference in confidence interval widths is about 0.011 AUROC units, which could be substantial relative to the often small improvements achieved during hyperparameter tuning. Hence, the resolving power analysis predicts that the AUROC will be better than the AUPRC for model search. But is it actually? The next section considers the challenge of empirically assessing the resolving power framework.
7 Empirical validation
Resolving power’s empirical importance crucially depends on the topography of the model search space. Its impact will be limited when search occurs in either “signal dominant” or “noise dominant” contexts. For instance, if the model evaluation points are spread across a steep region of the space then the model quality signal will overwhelm the noise for all metrics. In this case there will be consensus among metrics so it will not matter which is used for model selection. Conversely, some evaluation points may span an optimum that forms a flat region of the search space. If the flat portion is the entire search space or is sufficiently elevated (imagine an isolated butte) then again the choice of evaluation metric will not matter. But instead of unanimity we should now expect the metrics to agree only some of the time. [^7] If the metric estimates are independent across the optimum then the probability of consensus will be $(1/n)^{k-1}$ where $n$ is the number of evaluation points on the optimum and $k$ is the number of metrics.
The above concerns noted, we still want to test the hypothesis that higher resolving power leads to better model selection. An immediate challenge is that for actual data, such as with the diabetes readmissions example, we do not have access to the population. Yet with a sufficiently large dataset we can approximate drawing samples from a known population. This is the approach kohavi1995study uses to compare how different resampling methods affect classifier accuracy. Since the population characteristics of real data cannot be known, Kohavi uses the holdout method to estimate the “true” accuracies with relatively large test sets.
[…]
The AUROC had greater baseline estimated resolving power than the AUPRC on 499 of the 500 trials, meaning that the AUROC would be preferred for model search. For simplicity, the one aberrant trial is excluded from the subsequent analysis. The two metrics agreed on the best model on 281 out of the remaining 499 trials. Of the 218 disagreements, using the AUROC for model selection led to better AUROC test set performance on 146 of the trials and better AUPRC performance on 134 of the trials. That is, it was better on 67 and 61 percent of trials, higher than the predicted 50 percent if the two metrics were equally good for search.
Table 3: The top row gives the average performance for trials where ROC and PRC model choice disagrees. Disagreement occurred on 218 out of 499 total trials. The bottom row gives performance on the trials where the two metrics agree.
agree tune ROC, test ROC tune PRC, test ROC tune ROC, test PRC tune PRC, test PRC
FALSE 0.6205 0.6168 0.1495 0.1478
TRUE 0.6227 0.6227 0.1509 0.1509
[…]
For a given application the importance of resolving power will depend on the classification task, the candidate evaluation metrics, and the set of models being compared. Consequently, the observed effects of metric choice, such as those from this section, are liable to shrink or grow with any changes made to the model search space. This sensitivity to the problem context poses a special challenge for assessing the general empirical importance of resolving power.
@@ 8 Conclusion
Evaluation metrics form the contours of a model’s performance topography, so choosing the right metric is essential for successful navigation of this space. Resolving power is a framework for comparing threshold-free metrics. Central to the method is the specification of a class score sampling model that is used to both manipulate model quality and probe sampling variability. The quality dimension, which serves as the standard for comparison, is an ordered sequence of increasingly separated score distributions. A pivotal assumption is that movement on the quality dimension resembles the actual process of model improvement. This paper uses simple additive shifts to separate the class scores. Future work can test and refine this assumption by observing how risk scores evolve as algorithms learn from real-world data. Signal curves show how evaluation metrics respond to changes in classifier quality. Metric error variance is found with random draws from the sampling model. Resolving power is classifier-agnostic since it operates on risk scores that are downstream of a classification model. Binormal model simulation results provide general rules-of-thumb for when the AUROC will have stronger resolving power than the AUPRC. The empirical method allows researchers to use their data and an initial classifier of their choice to construct a sampling model for estimating resolving power.
If we imagine evaluation metrics occupying a 2-dimensional space with interpretability measured on the x-axis and resolving power on the y-axis, then a metric’s (intepretability, resolving power) coordinates describe its usefulness for model search. A (low, low) metric has no value. A (low, high) metric reliably maximizes an irrelevant or inscrutable target. A (high, low) metric fails to maximize the desired target. Of course (high, high) is best, but the challenge remains on how to create these types of metrics.
@@
All code and simulation data used in the paper are available on GitHub at https://github.com/colinbeam/resolving_power
@@ Appendix A Linear approximation method for resolving power
Inequality A.1 makes transparent the signal to noise comparison: The left side is the relative signal of the two metrics while the right side is the relative noise. If the inequality holds it means that the signal of the AUROC overwhelms its noise, giving it relatively greater resolving power.
Figure 8: Signal curve plotting the AUROC reference scale on the y-axis. A linear approximation of the signal function at the evaluation point is shown in red.